### v2

# arXiv:2602.19919v2[cs.CL] 27 Feb 2026

Abstract.
Financial market movements are often driven by discrete financial events conveyed through news, whose impacts are heterogeneous, abrupt, and difficult to capture under purely numerical prediction objectives. These limitations have motivated growing interest in using textual information as the primary source of trading signals in learning‑based systems. Two key challenges hinder existing approaches: (1) the absence of large-scale, event-centric datasets that jointly model news semantics and statistically grounded market reactions, and (2) the misalignment between language model reasoning and financially valid trading behavior under dynamic market conditions. To address these challenges, we propose Janus-Q, an end-to-end event-driven trading framework that elevates financial news events from auxiliary signals to primary decision units. Janus-Q unifies event-centric data construction and model optimization under a two-stage paradigm. Stage I focuses on event‑centric data construction, building a large‑scale financial news event dataset comprising 62,400 articles annotated with 10 fine‑grained event types, associated stocks, sentiment labels, and event‑driven cumulative abnormal return (CAR). Stage II performs decision‑oriented fine‑tuning, combining supervised learning with reinforcement learning guided by a Hierarchical Gated Reward Model (HGRM), which explicitly captures trade‑offs among multiple trading objectives. Extensive experiments demonstrate that Janus-Q achieves more consistent, interpretable, and profitable trading decisions than market indices and LLM baselines, improving the Sharpe Ratio by up to 102.0% while increasing direction accuracy by over 17.5% compared to the strongest competing strategies.
Event Trading, Reinforcement Learning, SFT
1. Introduction
Financial markets have long been studied through the lens of time-series modeling, where asset prices or returns are predicted directly from historical numerical signals such as prices, volumes, and technical indicators (Zhou et al., 2025b; Zhang et al., 2025b). Despite decades of methodological advances, purely numerical forecasting remains fundamentally challenging due to severe noise, non-stationarity, and frequent regime shifts in real-world markets (Fama, 1970; Lo, 2004). These limitations have motivated growing interest in incorporating alternative information sources, including financial news, corporate disclosures, and social media, into learning-based trading systems (Zhou et al., 2025c; Wang et al., 2026; Zhou et al., 2024).
Figure 1. Comparison of traditional time-series–driven trading and event-driven trading strategies.
However, a key characteristic of real financial markets is often overlooked: asset price movements are rarely driven by smooth temporal dynamics alone. Instead, they are frequently precipitated by discrete and interpretable events, such as earnings announcements, mergers and acquisitions, or risk disclosures, which can abruptly shift investor expectations and asset valuations (MacKinlay, 1997a; Thompson, 1995). These events constitute the primary mechanism through which new information is incorporated into prices. Crucially, different event categories induce highly heterogeneous market responses in terms of direction, magnitude, and temporal persistence. Treating such structurally distinct events as homogeneous signals inevitably obscures their economic meaning and limits decision quality.
Although most market-moving events are communicated through unstructured financial news, existing learning-based trading systems have yet to fully exploit event-level information. The dominant paradigm embeds textual news and fuses it with historical price sequences under numerical forecasting objectives (Kong et al., 2025a; Zhang et al., 2025c). Under this formulation, textual information is typically treated as an auxiliary modality, rather than as a primary driver of trading decisions (Li et al., 2024b; Dong et al., 2024; Saqur et al., 2024). As a result, learned policies are often dominated by recent price dynamics, even when those dynamics contradict the semantic implications of newly released events. Figure 1 illustrates this structural mismatch between time-series–driven and event-driven trading paradigms. This limitation stems primarily from the lack of structured supervision at the event level. While prior work has explored sentiment labels or coarse market reactions (Tetlock, 2007a; Wang et al., 2025a; Li et al., 2025), two fundamental challenges remain largely unaddressed.
Challenge 1 (Ch1): Lack of Event–Market Granularity. Existing financial datasets fail to jointly model what event occurred, which assets were affected and with what semantic polarity, how the market responded over time, and the coverage of event domains. Without fine-grained event categories and statistically grounded post-event outcomes, models struggle to distinguish economically meaningful news from noise or to reason about heterogeneous market responses across event types (Chen et al., 2025; Sinha et al., 2022; Han et al., 2022). As shown in Table 1, most existing resources either focus on narrow domains (e.g., risk or macro events) or provide coarse labels without jointly modeling sentiment and impact magnitude (e.g., CAR). General-domain datasets suitable for event-driven trading remain particularly scarce. Moreover, approaches that infer event impact directly from raw price sequences are heavily confounded by market-wide and firm-specific factors, making it difficult to isolate causal event effects or learn heterogeneous event-to-market mappings.
Challenge 2 (Ch2): Misalignment Between Semantic Reasoning and Market Reality. Beyond data limitations, aligning language model reasoning with empirical market behavior remains nontrivial. Although large language models (LLM) can generate fluent and plausible interpretations of financial news, their semantic judgments are not inherently grounded in realized market outcomes (Zhang et al., 2024; Zhou et al., 2025a). In practice, semantic polarity does not map linearly to price movements: apparently positive announcements may trigger corrections due to priced-in expectations, while negative news can be absorbed by the market and followed by stabilization or rebound. Consequently, purely supervised learning risks capturing superficial correlations, whereas profit-driven optimization alone often induces spurious strategies that exploit short-term noise (Liu et al., 2024).
To address these challenges, we propose Janus-Q, a two-stage event-driven trading framework that elevates financial news events from auxiliary features to primary decision units. First, to tackle (Ch1), we construct a large-scale financial event dataset comprising 62,400 expert-annotated news articles, each labeled with fine-grained event types, associated stocks, semantic polarity, and cumulative abnormal returns (CAR), which capture event-induced abnormal returns by aggregating deviations from expected, event-free price dynamics over statistically defined event windows.
Building on this dataset, we develop a decision‑oriented training paradigm that directly maps financial news events to executable trading actions. As illustrated in Figure 2, the framework adopts a multi‑step optimization paradigm. Supervised fine‑tuning first establishes a reasoning‑aware mapping from event descriptions to expected CAR, effectively integrating textual semantics, market signals, and firm‑specific features. Subsequently, a Hierarchical Gated Reward Model (HGRM) is introduced for reinforcement fine‑tuning. This component explicitly addresses the semantic–market gap (Ch2) by decomposing the trading reward into interpretable components reflecting event‑type consistency, directional accuracy, and return‑magnitude reliability. The hierarchical structure further acts as a form of semantic regularization, constraining the learned policy to realize profits through financially grounded reasoning rather than spurious reward exploitation, while maintaining sensitivity to risk and transaction costs.
Extensive experiments demonstrate that Janus-Q consistently outperforms market indices, time-series–oriented models, and both vanilla and financial LLM baselines. On average, Janus-Q improves direction accuracy by 17.5% and event-type accuracy by 18.2% over the strongest competing methods. Ablation studies further confirm the critical role of HGRM in achieving these gains. The principal contributions of this paper are as follows:
  • We construct a large-scale dataset of 62,400 financial news articles manually annotated with 10 event types, associated stocks, semantic labels, and event-driven abnormal returns, forming a unified benchmark for event-level market impact analysis.
  • We propose Janus-Q, the first end-to-end event-driven trading framework that directly maps financial news events to trading decisions, unifying event interpretation and market response learning via HGRM-guided optimization.
  • We evaluate Janus‑Q through extensive experiments and find that it delivers strong and consistent trading performance, improving the Sharpe Ratio by 102.0% and direction accuracy by 5.4% compared with the runner-up strategy, while maintaining a comparable maximum drawdown. These results confirm the effective alignment between event understanding and trading decisions.
Table 1. Comparison of key characteristics of financial event datasets, including event typing, sentiment annotation, market impact magnitude, domain scope, event-type granularity, dataset scale, and language coverage.
Dataset Event Typing Open Source Sentiment Impact Magnitude (CAR) Domain # Event Types Size Language
ChFinAnn (Zheng et al., 2019) Risky 5 32,040 CN
DuEE-Fin (Han et al., 2022) General 13 15,000 CN
FNSPID (Dong et al., 2024) / / 15,700,000 EN
SEntFiN (Sinha et al., 2022) / / 10,753 EN
CAMEF (Zhang et al., 2025c) Macro 7 2,596 EN
CSMAR (Wang et al., 2025b) General 13 / CN
DocFEE (Chen et al., 2025) Risky 9 19,044 CN
Ours General 10 62,400 CN
2. Related Work
2.1. Event-Driven Modeling
Early studies on event-driven market analysis originate from econometric and statistical finance, where discrete corporate or macroeconomic events are treated as exogenous shocks and analyzed through abnormal return dynamics around predefined event windows. The classical event study methodology formalized this process by estimating expected returns from factor models and attributing deviations to the information content of events, establishing a standard toolkit for measuring market reactions (MacKinlay, 1997b). As textual disclosures and news became increasingly accessible, subsequent research incorporated qualitative information into event analysis, showing that media tone and sentiment convey economically meaningful signals that explain and predict short-horizon market movements beyond purely price-based features (Tetlock, 2007b). Building on this line, machine learning (ML) and deep learning (DL) approaches began to integrate textual information with market data, shifting from handcrafted sentiment indicators toward learning event semantics directly from text. Ding et al. (Ding et al., 2015) extract structured events from news and learn dense event representations to model both short-term and long-term effects on stock movements, with subsequent work further enhancing event representations by incorporating richer contextual and temporal signals (Ding et al., 2016). More recently, Wang et al. (Wang et al., 2025b) introduce StockMem, an event-reflection memory framework that organizes news into structured events and leverages their temporal evolution to retrieve analogous historical scenarios, supporting more explainable stock movement prediction. However, despite richer event representations, existing pipelines remain prediction‑oriented and treat events primarily as inputs, without explicitly modeling event‑level properties such as historical context or relative importance to support decision‑oriented trading objectives. Moreover, the field still lacks a comprehensive, standardized dataset for trading. Following this direction, we propose an event‑centric dataset that integrates event categorization, sentiment annotation, and CAR‑based evaluation of event impacts.
2.2. LLM for Trading
Recent studies have explored LLMs for trading-related tasks, motivated by their ability to process unstructured financial texts and to provide explicit reasoning through chain-of-thought generation (Xie et al., 2024a; Yu et al., 2024). Unlike classical ML or DL models that operate as black-box predictors relying on past numerical data and struggling with abrupt market shifts, LLMs can articulate intermediate causal rationales when interpreting complex market events and respond adaptively to sudden market shocks (Tatsat and Shater, 2025; Cao et al., 2025). Yang et al.(Yang et al., 2025b) adapt instruction‑tuned LLM for financial analysis and trading, showing that they effectively extract market‑relevant semantics from financial texts. Zhang et al.(Zhang et al., 2024) propose FinAgent, an LLM‑based financial agent that interacts with market environments and external tools to support trading and portfolio management. Xiao et al. (Xiao et al., 2024) study multi-agent trading frameworks powered by LLM, where language models coordinate decision-making through communication and tool usage in simulated markets. More recently, Xiao et al. (Xiao et al., 2025) introduce Trading-R1, which applies reinforcement learning to enhance LLM reasoning for trading decisions, highlighting the potential of RL-based fine-tuning for aligning language model outputs with trading actions. In parallel, Li et al. (Lin et al., 2025) propose RETuning, a framework that improves stock movement prediction by refining inference-time reasoning with reflective and evidence-based analysis over rich textual inputs. Despite these advances, current LLM-based trading systems face two complementary limitations. First, trading decisions are often treated as opaque actions primarily driven by price or volume signals, with textual information only weakly integrated into the decision process. Second, reinforcement learning–based methods typically rely on heuristic, linearly additive reward designs in which competing objectives may offset one another, limiting their ability to model economically meaningful trade-offs. Motivated by these limitations, we propose a Hierarchical-Gated Reward Model that aligns event-level semantic reasoning with market-grounded trading outcomes.
Figure 2. Overview of the proposed two-stage event-driven trading framework. Stage I focuses on event-centric data construction with market-grounded supervision, while Stage II performs decision-oriented fine-tuning via supervised and reinforcement learning.
3. Methodology
This section describes the why and how of event-driven trading. As illustrated in Figure 2, our framework is divided into two core paradigms: Stage I, which focuses on data construction, and Stage II, which focuses on model training.
3.1. Event Centric Data Construction
Why: The natural question guiding the collection of this dataset is: how will this event affect the associated asset in terms of direction and magnitude, and how does its impact differ from that of other event types? How: To address this, we construct an event-centric dataset that captures both the market impact and the semantic interpretation of financial news events. Specifically, motivated by classical event study methodology (MacKinlay, 1997a), we first quantify impact magnitude through Event-to-CAR modeling, and then associate each event with category labels and structured semantic annotations to support event-level supervision. Event type annotations are conducted by a panel of six domain professionals, including fund researchers and securities analysts, ensuring consistency, financial relevance, and reliability.
3.1.1. Event-to-CAR Modeling
Given a news event for stock $i$ occurring at event time $t_{0}$, we quantify its market impact using the CAR computed over a predefined event window. Let $r_{i,t}$ denote the return of stock $i$ on trading day $t$. We define two time intervals: an estimation window $\mathcal{T}_{\mathrm{est}}=(T_{0},\,T_{1}]$ precedes the event and is used to estimate event-free normal returns, and an event window $\mathcal{T}_{\mathrm{evt}}=(T_{1},\,T_{2}]$ captures abnormal price reactions associated with the event, including potential pre-event leakage and delayed market adjustment. Here $T_{0},T_{1},$ and $T_{2}$ specify the temporal boundaries of the estimation and event windows relative to $t_{0}$. The timing structure is illustrated in Figure 5.
$\bullet$ Market Model (MR) To remove broad market movements, we estimate the expected return of each stock $i$ using the market model over the estimation window $\mathcal{T}_{\mathrm{est}}$:
(1) $r_{i,t}=\alpha_{i}+\beta_{i}r_{m(i),t}+\varepsilon_{i,t},\quad t\in\mathcal{T}_{\mathrm{est}}.$
where $r_{m(i),t}$ is the return of the benchmark market index selected for stock $i$. We choose the benchmark index $m(i)$ according to the market-cap segment of stock $i$ (e.g., large-/mid-/small-cap) so that $r_{m(i),t}$ matches its investable universe.
Given the parameters $\hat{\alpha}_{i}$ and $\hat{\beta}_{i}$ estimated from the estimation window $\mathcal{T}_{\mathrm{est}}$ via an ordinary least squares fit of the market model (Kolari and Pynnönen, 2010), we evaluate the fitted model on the event window to obtain abnormal returns (AR) as
(2) $AR^{\mathrm{MR}}_{i,t}=r_{i,t}-\bigl(\hat{\alpha}_{i}+\hat{\beta}_{i}r_{m(i),t}\bigr),\quad t\in\mathcal{T}_{\mathrm{evt}}.$
$\bullet$ Risk Model (RM) Neutralization Market-adjusted returns may still contain systematic style/industry effects. We further neutralize $AR^{\mathrm{MR}}_{i,t}$ using a multi-factor risk model on $\mathcal{T}_{\mathrm{evt}}$:
(3) $AR^{\mathrm{MR}}_{i,t}=\gamma_{i}+\mathbf{x}_{i,t}^{\top}\boldsymbol{\lambda}_{t}+u_{i,t},\quad t\in\mathcal{T}_{\mathrm{evt}}.$
where $\mathbf{x}_{i,t}$ denotes the Barra factor exposures of stock $i$ (e.g., style and industry exposures under risk model), $\boldsymbol{\lambda}_{t}$ are the corresponding factor returns/premia at time $t$, $\gamma_{i}$ is a stock-specific intercept capturing persistent effects not explained by the factors, and $u_{i,t}$ is the idiosyncratic component.
We apply the estimated factor premia to obtain the factor-neutral abnormal return within the event window:
(4) $AR_{i,t}=AR^{\mathrm{MR}}_{i,t}-\mathbf{x}_{i,t}^{\top}\widehat{\boldsymbol{\lambda}}_{t},\quad t\in\mathcal{T}_{\mathrm{evt}}.$
Details of the risk model specification and factor definitions are provided in Appendix A.3.
Cumulative Abnormal Return (CAR). We then aggregate abnormal returns over the event window to obtain the cumulative abnormal return:
(5) $CAR_{i}=\sum_{t\in\mathcal{T}_{\mathrm{evt}}}AR_{i,t}.$
The resulting CAR summarizes the total abnormal price movement attributable to the event, capturing both potential pre-event information leakage and delayed post-event market adjustment.
3.1.2. Event Taxonomy and Sentiment Annotation
The cumulative abnormal return (CAR) defined above serves as the impact magnitude of an event, quantifying its realized economic effect on asset prices. To complement this market-grounded signal, we further associate each event with structured semantic labels.
Formally, each event is represented by a ground-truth label $y=\{c,d,s,e\}$, where $c$ denotes the realized CAR and $e\in\mathcal{E}$ is the annotated event type. In our formulation, the direction $d$ (positive, negative, or neutral) and the trading strength $s$ (strong or weak) are deterministically derived from $c$: the direction is obtained via the sign function $d=\mathrm{sign}(c)$, while the trading strength is determined by a predefined threshold $\tau$ that specifies whether the magnitude of the market impact is sufficient to justify the execution of the trade. Together, $d$ and $s$ serve as semantic that characterizes how the market reacts to an event, complementing the quantitative impact magnitude $c$. The detailed taxonomy of event types, the corresponding annotation criteria, and the dataset format are provided in Appendix A.1.2.
3.2. Decision-Oriented Finetuning
Why: (1) Purely supervised finetuning cannot ensure that model reasoning translates into market-consistent decisions. (2) Existing reinforcement learning–based methods often rely on heuristic, additively designed rewards, where competing objectives may offset each other, limiting their ability to model economically meaningful trade‑offs. How: Building on the event-centric dataset, we perform decision-oriented fine-tuning to align model reasoning with executable trading actions. This stage follows a multi-step training paradigm. We first apply supervised fine-tuning (SFT) to stabilize structured event reasoning, and then perform reinforcement fine-tuning using Group Relative Policy Optimization (GRPO) to directly optimize trading decisions. To ensure that reinforcement learning is guided by economically meaningful signals, we design a structured reward mechanism, detailed below.
3.2.1. Hierarchical-Gated Reward Modeling
To bridge semantic event understanding with executable trading actions, we propose a Hierarchical-Gated Reward Model (HGRM) that provides structured supervision for reinforcement-based fine-tuning. For a given stock $i$ at event time $t_{0}$, the model generates a response that is interpreted as a composite prediction $\hat{y}=\{\hat{c},\hat{d},\hat{s},\hat{e}\}$, where $\hat{c}$ denotes the predicted cumulative abnormal return (CAR), $\hat{d}$ represents the predicted direction (positive, negative, neutral), $\hat{s}$ indicates whether a trade should be executed (strong or weak). and $\hat{e}$ corresponds to the inferred event type.
[…]
3.2.2. Hard gate: Direction correctness.
We first enforce a hard direction gate $g_{\mathrm{dir}}\in\{0,1\}$ to prevent spurious profits under incorrect market polarity. The direction score is defined as
(6) $s_{\mathrm{dir}}(\hat{d},d)=\begin{cases}1,&\hat{d}=d,\\ -\lambda_{\mathrm{dir}},&\hat{d}=-d,\\ 0,&\text{otherwise},\end{cases}$
where $\lambda_{\mathrm{dir}}>1$ assigns a stronger penalty to opposite directions than to ambiguous cases involving neutral. When $s_{\mathrm{dir}}<0$, we set $g_{\mathrm{dir}}=0$ to block all subsequent reward contributions and avoid rewarding trades with incorrect directional alignment.
3.2.3. Soft gate: Event-type consistency.
We introduce an event-type soft gate that discounts rewards when the predicted event category is incorrect:
(7) $s_{\mathrm{evt}}(\hat{e},e)=\begin{cases}1,&\hat{e}=e,\\ -\lambda_{\mathrm{evt}},&\hat{e}\neq e,\\ -\lambda_{\mathrm{miss}},&\hat{e}=\emptyset,\end{cases}\qquad m_{\mathrm{evt}}=\begin{cases}1,&\hat{e}=e\\ \alpha,&\text{otherwise},\end{cases}$
Here, $s_{\mathrm{evt}}(\hat{e},e)$ provides a signed supervision signal for event-type correctness, while $m_{\mathrm{evt}}$ acts as a multiplicative discount on trading rewards. The penalty parameters $\lambda_{\mathrm{evt}}>0$ and $\lambda_{\mathrm{miss}}>0$ control the strength of penalization for incorrect and missing event-type predictions (i.e., $\hat{e}$$\emptyset$), respectively. The discount factor $\alpha\in(0,1)$ reduces the contribution of profit-based rewards when the predicted event type does not match the ground truth, encouraging consistency between the understanding of the event and the trading outcomes.
3.2.4. Trading reward: cost-aware PnL with strength regularization.
To reflect economically meaningful outcomes, we define a cost-aware single-event trading payoff based on the realized CAR $c$:
(8) $\mathrm{PnL}(\hat{d},c)=\begin{cases}c-\kappa,&\hat{d}=\texttt{positive},\\ -c-\kappa,&\hat{d}=\texttt{negative},\\ 0,&\hat{d}=\texttt{neutral},\end{cases}$
where $\kappa$ denotes the transaction cost. The payoff is activated only when the predicted trading strength is $\hat{s}=\texttt{strong}$, scaled by the event-type discount factor $m_{\mathrm{evt}}$ and the realized profit and loss $\mathrm{PnL}(\hat{d},c)$, and set to zero otherwise. To stabilize reinforcement learning, the resulting reward is clipped to a symmetric range $\rho,\rho]$, where $\rho>0$ is a predefined bound. We denote the resulting clipped profit-based reward as $r_{\mathrm{pnl}}=\mathrm{clip}(m_{\mathrm{evt}}\cdot\mathrm{PnL}(\hat{d},c),-\rho,\rho)$.
In addition, we regularize the predicted trading strength to discourage degenerate strategies. False-positive decisions (predicting strong when no trade is warranted) and false-negative decisions (predicting weak when profitable opportunities exist) are penalized with asymmetric costs, preventing the model from collapsing to always-trade or never-trade behaviors.
3.2.5. Magnitude shaping and process reward.
To provide fine-grained supervision beyond direction, we add a magnitude shaping term when the direction is not blocked:
(9) $r_{\mathrm{mag}}=\exp\!\left(-\frac{|\hat{c}-c|}{\sigma}\right),$
where $\sigma$ controls the tolerance scale. We also include a lightweight process reward $r_{\mathrm{proc}}$ that evaluates the presence of required reasoning sections and penalizes overly long responses and excessive self-questioning.
3.2.6. Overall reward.
We compose the final HGRM reward hierarchically, where higher-level gates control whether lower-level rewards are activated:
(10) $R$ $=w_{\mathrm{dir}}\,s_{\mathrm{dir}}$
$\quad+g_{\mathrm{dir}}\Big(w_{\mathrm{evt}}\,s_{\mathrm{evt}}+w_{\mathrm{pnl}}\,r_{\mathrm{pnl}}+w_{\mathrm{mag}}\,r_{\mathrm{mag}}+w_{\mathrm{proc}}\,r_{\mathrm{proc}}\Big).$
where $g_{\mathrm{dir}}\in\{0,1\}$ blocks lower-level rewards when the predicted direction is opposite to the realized movement (i.e., $s_{\mathrm{dir}}<0$). The Algorithm 1 formalizes the hierarchical gated reward modeling procedure. The corresponding training configurations and implementation details are provided in Appendix A.1.1.
4. Experiment Results
4.1. Experiment Setup
Given a universe of stocks $\mathcal{S}$, for any stock $s\in\mathcal{S}$ on a trading day $t$, we evaluate a long–short trading strategy driven by Janus-Q. The trading process is defined as follows. (1) Signal Generation. Janus-Q analyzes all news events released between the market open at 9:309{:}30 on day $t$ and the market open at 9:309{:}30 on day t+1t{+}1, and produces a directional signal $\gamma_{s,t}\in\{\text{Long},\text{Short},\text{Hold}\}$ for each stock $s$. (2) Event-weighted Signal Aggregation. On each trading day $t$, news signals are aggregated at the portfolio level using event-type weights $w_{k}$ estimated from historical post-event abnormal returns. The daily budget is allocated across event types and evenly split within each type using only information available before day t+1t{+}1 market open. (3) Entry Rule. If the aggregated signal $\gamma_{s,t}$ indicates a long (short) position, we initiate a long (short) trade at the opening price $o_{s,t+1}$ on day t+1t{+}1. (4) Exit Rule. The position is held until the last trading day within the subsequent two trading days, denoted as $\tau(t)$, at which point the stock is sold at the closing price $c_{s,\tau(t)}$.
4.2. Evaluation Objective
We conduct a comprehensive evaluation to assess the effectiveness of Janus-Q in event-driven trading. Our evaluation is designed to examine both overall trading performance and the contribution of key design components, with a particular focus on decision alignment, reward modeling, and human-consistent understanding. To this end, we structure our experiments around the following research questions (RQs):
$\bullet$ RQ1: Does Janus-Q consistently outperform market indices and competitive model baselines in terms of trading performance and decision accuracy?
$\bullet$ RQ2: How does each core component of Janus-Q contribute to its overall trading performance?
$\bullet$ RQ3: How effective are diversified reward objectives in improving learning stability and decision-oriented trading performance?
$\bullet$ RQ4: To what extent does Janus-Q align with human judgments in interpreting financial events?
4.3. Dataset & Metrics
We evaluate our model on a multi-source dataset that integrates both textual and financial information. Raw news articles are collected from the Datayes platform [^1] https://www.datayes.com, spanning the period from January 1, 2023 to January 25, 2025. Corresponding stock price data for backtesting are retrieved from Tushare [^2] https://tushare.pro, covering the period from January 1, 2023 to February 6, 2025. All data are chronologically split into a 4:4:1:1 ratio for historical statistics, training, validation, and testing. Additionally, we incorporate corporate profiles, including industry category and market share—sourced from the Wind platform [^3] https://www.wind.com.cn to support semantic enrichment.
To evaluate the performance of our model, we adopt six metrics: Mean Absolute Error (MAE), Root Mean Square Error (RMSE), Direction Accuracy (DA), Event Type Accuracy (ETA), Sharpe Ratio (SR), and Maximum Drawdown (MDD). Together, these metrics provide a comprehensive view of both predictive quality and practical investment performance of our model.
4.4. Baseline
Our proposed approach is evaluated against four categories of baselines:
$\bullet$ Market Indices. Market indices serve as standard passive investment benchmarks. We report results on several representative indices, including the CSI 300, CSI 500, and CSI 1000, which collectively cover large-, mid-, and small-cap segments of the Chinese equity market.
$\bullet$ Time-Series-Oriented LLM. We include a set of LLM specifically designed for modeling temporal and sequential data, which adapt LLM architectures to time-series forecasting tasks. Representative models in this category include Time-MQA (Kong et al., 2025b), ChatTS-14B (Xie et al., 2024b), and TimeMaster (Zhang et al., 2025a).
$\bullet$ Financial Domain-Specific LLM. We evaluate several open-source language models pre-trained or fine-tuned on financial corpora, including FinMA (Xie et al., 2023), DISC-FinLLM (Chen et al., 2023), and Stock-Chain (Li et al., 2024a). These models are designed to capture financial semantics and domain-specific patterns, enabling direct application to market analysis and trading-related evaluation.
$\bullet$ General-Purpose LLM. We further include widely used general-purpose language models, including QwQ-32B (Yang et al., 2025a), Claude-3-Haiku (Rahman et al., 2024), GPT-4o-mini (Hurst et al., 2024), DeepSeek-v3.1-nex-n1 (Cai et al., 2025), Grok-3-mini-beta (Jiang et al., 2025), Qwen2.5-7B (Team et al., 2024), and Gemini-2.5-flash (Comanici et al., 2025), which are not specifically tuned for financial tasks.
Figure 3. Cumulative returns of each baseline strategy on the financial news dataset from November 12, 2024 to February 6, 2025. The figure shows the net asset value (NAV) curves over the backtesting period.
Table 2. Performance comparison across Market Indices, Vanilla LLM, Time-aware LLM, and Financial LLM.
Model MAE↓ RMSE↓ DA↑ ETA↑ SR↑ MDD↓
Market Indices CSI 300 -1.8027 0.0945
CSI 500 -1.6074 0.1423
CSI 1000 -0.1036 0.1456
TimeLLM Time-MQA 0.0427 0.2651 0.4732 0.5958 -4.2184 0.1589
ChatTS-14B 0.0932 0.3271 0.4995 0.6223 -4.2827 0.1675
TimeMaster 0.0472 0.1015 0.4472 0.6263 -5.9237 0.1981
FinLLM Stock-Chain 0.1352 0.3226 0.4510 0.4364 -5.3581 0.1800
FinMA 0.0947 0.2531 0.4550 0.5714 -6.9668 0.2358
DISC-FinLLM 0.0707 0.1672 0.4286 0.5608 -4.6307 0.2139
Vanilla LLM QwQ-32B 0.0437 0.0839 0.4674 0.7080 0.6481 0.0979
Claude-3-Haiku 0.0384 0.0557 0.4547 0.4884 -5.4698 0.1395
GPT-4o-mini 0.0387 0.0558 0.4700 0.7508 -3.8759 0.1060
DeepSeek-v3.1-nex-n1 0.0541 0.1214 0.4365 0.6811 0.3710 0.1300
Grok-3-mini-beta 0.0369 0.0559 0.4795 0.5417 -2.2698 0.0855
Qwen2.5-7B 0.0377 0.0551 0.4114 0.4879 -3.9154 0.1177
Gemini-2.5-flash 0.0534 0.0730 0.4276 0.6797 -2.0899 0.0805
Finetuning LLM Janus-Q (Ours) 0.0349 0.0541 0.5869 0.8009 1.3088 0.1196
4.5. Experimental Results
4.5.1. Baseline Comparison (RQ1)
We compare Janus-Q against 16 baseline methods across six general evaluation metrics. Figure 3 depicts a market regime characterized by a correction phase, followed by a broad pullback and consolidation, reflecting the waning influence of favorable Chinese fiscal policy signals and subsequent index-level corrections. During this period, all three benchmark indices fail to sustain positive risk-adjusted returns, with Sharpe Ratios remaining negative throughout. Most time-aware models and financial LLM exhibit similar patterns, either enduring persistent drawdowns or demonstrating weak, oscillatory NAV trajectories that closely mirror the underlying market dynamics. Among general-purpose LLM, only QwQ-32B and DeepSeek-v3.1-nex-n1 achieve positive Sharpe Ratios; however, their NAV curves remain highly volatile and lack sustained upward momentum. In contrast, Janus-Q exhibits a distinctly different trajectory. It effectively captures the sharp upswing observed in late December and maintains a stable and consistent growth trend thereafter. By the end of the evaluation window, Janus-Q achieves the highest cumulative return among all compared methods, highlighting its superior ability to adapt to shifting market conditions and translate predictive precision into tangible trading gains.
The quantitative results in Table 2 corroborate these findings. Janus-Q consistently outperforms all baselines on both decision-level and trading-level metrics. It achieves the lowest CAR prediction error, reducing MAE by 18.3% relative to the best time-aware model and by 50.6% over the strongest financial LLM baseline; even against the best vanilla LLM, Janus-Q still lowers error by 5.4%. Its direction accuracy further exceeds the best time-aware, financial, and vanilla LLM baselines by 17.5%, 29.0%, and 22.4%, respectively. Notably, domain-specific FinLLM underperform most vanilla LLM, likely due to task–domain mismatch during finetuning and architectural or context-length constraints that hinder effective modeling of long, complex news narratives. Meanwhile, time-aware models remain inferior in backtesting performance despite explicitly incorporating textual features. This indicates that merely injecting textual cues as auxiliary inputs is insufficient for robust trading, reinforcing the value of directly optimizing the event-to-decision mapping. These decision-level gains translate into significantly better trading outcomes. Janus-Q attains a Sharpe Ratio of 1.3088, more than doubling that of the runner-up QwQ-32B and yielding a relative improvement above 102.0%. In contrast, both time-aware and financial LLM baselines exhibit negative Sharpe Ratios over the same test horizon. Although some baselines achieve marginally lower drawdowns, this comes at the cost of substantially reduced returns, whereas Janus-Q attains a more favorable balance between profitability and stability with superior risk-adjusted performance.
4.6. Ablation Study
4.6.1. Effectiveness of Each Component (RQ2)
Table 3 reports the ablation study evaluating the contribution of each main component in Janus-Q. Removing any component consistently degrades performance, confirming that the framework’s effectiveness arises from the synergy among its core modules. Among all variants, eliminating supervised fine-tuning leads to the most pronounced deterioration, with direction accuracy declining by over 14% and trading performance shifting from a robustly positive regime to a distinctly negative Sharpe Ratio. This finding underscores its fundamental role in establishing a reliable decision-making foundation. In comparison, removing reinforcement optimization causes a moderate yet consistent decline of approximately 13% in Sharpe Ratio, suggesting that reinforcement learning primarily fine-tunes and enhances an already acquired policy rather than replacing supervised learning. Ablating CAR supervision or company-level information results in smaller but still noticeable reductions in both predictive accuracy and profitability, demonstrating that market impact cues and firm-specific context make meaningful contributions to effective event-driven trading.
Table 3. Ablation study of Janus-Q. Each variant removes a single source-level or structural component from the full model.
Methods MAE DA ETA SR
w/o CAR 0.0387 0.5261 0.7796 0.8690
w/o Co. Info 0.0353 0.5464 0.7790 0.9608
w/o SFT 0.0381 0.4429 0.6771 -5.2848
w/o GRPO 0.0355 0.5459 0.7881 1.1330
Janus-Q (Full) 0.0349 0.5869 0.8009 1.3088
4.6.2. Effectiveness of Diversified Reward Objectives (RQ3)
Table 4. Ablation study of diversified reward objectives in HGRM. Each variant removes a single reward component from the full model.
Methods MAE DA ETA SR
w/o Event Type 0.0345 0.5819 0.7817 1.2107
w/o Direction 0.0357 0.5589 0.7896 1.2788
w/o Magnitude 0.0347 0.5653 0.7843 1.1558
w/o PNL 0.0354 0.5799 0.7868 1.1953
HGRM (Full) 0.0349 0.5869 0.8009 1.3088
Table 4 examines the impact of removing individual reward objectives from the full HGRM. Overall, all ablated variants underperform the full model, indicating that diversified reward objectives play complementary roles in learning stable and effective trading policies. Removing the direction-related objective leads to the most pronounced drop in decision quality, with direction accuracy decreasing by over 4.8%, while the impact on trading performance remains comparatively moderate, likely because the model still captures a large fraction of high-impact signals. In contrast, removing event-type supervision slightly reduces prediction error and leads to a modest decline in trading performance, as the absence of event-type signals affects position weighting and weakens the allocation of capital across heterogeneous events. A detailed analysis of this effect is provided in Appendix, as illustrated in Figure 7 and Figure 8. Eliminating magnitude or profit-and-loss objectives results in more pronounced performance degradation, reducing the Sharpe Ratio by approximately 11.7% and 8.7%, respectively. In comparison, the full HGRM achieves the highest decision accuracy and the strongest trading performance, demonstrating that jointly optimizing diversified reward objectives is critical for stable learning and decision-oriented event-driven trading.
4.7. Case Study
4.7.1. Human Alignment in Event Interpretation (RQ4)
Figure 4 reports a direct comparison among Janus-Q, strong LLM baselines, and human evaluations in event interpretation. The evaluation is conducted on 200 randomly selected samples from the test dataset. To obtain a representative measure of human judgment, we recruit three participants with varying levels of financial expertise, including a finance major student, a securities analyst, and a CFA certificant. This design captures a spectrum of professional backgrounds and mitigates potential bias from individual knowledge gaps. Their judgments are aggregated through majority voting to form a balanced human consensus benchmark. For direction prediction, Janus-Q demonstrates substantial agreement with both humans and other models, achieving tie rates between 40.5% and 52.0%, win rates of up to 37.8% against DeepSeek-v3.1-nex-n1, and loss rates not exceeding 25.0%. The alignment becomes even more pronounced in event-type understanding, where tie cases dominate across comparisons, accounting for 74.0% to 83.0%, and loss rates remain below 5.0% for all counterparts, including human judges. In addition, Janus-Q secures non-trivial win rates ranging from 14.0% to 21.0%, suggesting that its deviations from human judgment are infrequent and occur in a controlled manner rather than reflecting systematic semantic discrepancies. Overall, these results show that Janus-Q demonstrates a higher level of accuracy in event interpretation compared to the average performance of human evaluators, as variations in human judgment may arise from the subjective nature of the evaluation. Janus-Q also retains the flexibility required for decision-oriented and context-sensitive trading behavior.
Figure 4. Visualization of comparative evaluation results between humans and models.
5. Conclusion
This paper investigates event‑driven trading as an alternative to conventional time‑series‑centric formulations and demonstrates that explicitly modeling financial news events as primary decision units leads to more reliable and interpretable trading behavior. We propose Janus‑Q, a two‑stage event‑driven trading framework that directly maps financial events to trading decisions. The framework first constructs a large‑scale event‑centric dataset that serves as a unified benchmark for event‑level market impact analysis by linking fine‑grained event semantics with empirically grounded market reactions. Building on this foundation, Janus‑Q adopts a multi‑step training paradigm that combines supervised reasoning alignment with reinforcement fine‑tuning under a hierarchical gated reward scheme. Extensive experiments show that Janus‑Q consistently surpasses both market indices and strong large‑language‑model baselines in decision accuracy and trading performance, highlighting the importance of aligning language model reasoning with financially meaningful objectives. In future work, we plan to extend this framework to support finer‑grained event structures, richer multimodal inputs, and cross-market backtesting, further advancing event‑driven learning for real‑world financial trading.
Appendix A Appendix
Table 5. SFT training hyperparameters.
Hyperparameter Value / Setting
Data Configuration
Max sequence length 1400
Dataset shuffling True
Training Configuration
Epochs $\{1,2,3\}$
Train micro-batch size per device 8
Gradient accumulation steps 4
Total effective batch size $8\times 4\times N_{\mathrm{GPU}}$
Precision BF16
Gradient checkpointing True
Evaluation steps 100
Logging steps 10
Optimizer & LR Scheduler
Optimizer torch.optim.AdamW
Learning rate $5\times 10^{-6}$
Warmup ratio 0.2
Weight decay 0.01
AdamW betas 0.9, 0.999
LoRA Configuration
LoRA rank $r$ 8
LoRA alpha $\alpha$ 16
LoRA dropout 0.1
Task type CAUSAL_LM
Table 6. GRPO training hyperparameters.
Hyperparameter Value / Range
Training Configuration
Epochs $\{1,2,3\}$
Batch size $\{1,2,4\}$
Learning rate $\{1\mathrm{e}{-6},2\mathrm{e}{-6},5\mathrm{e}{-6}\}$
Max grad norm $\{0.5,1.0\}$
Ref Configuration
PPO epochs $\{1,2\}$
Samples per prompt $\{2,4,8\}$
Initial KL coefficient $\{0.05,0.1,0.2\}$
Target KL $\{0.01,0.02\}$
Clip range $\{0.1,0.2\}$
Normalize advantage by std {True, False}
Data Configuration
Max prompt length $\{3000,4196\}$
Max response length $\{4196,8192\}$
Data Shuffling {True, False}
System Configuration
Number of GPUs $\{1,2,4,8\}$
Number of nodes $\{1,2\}$
vLLM inference {True, False}
DeepSpeed ZeRO stage {None, ZeRO-2, ZeRO-3}
Mixed precision {FP16, BF16}
A.1. Hyperparameters & Datasets & Metrics
A.1.1. Training details.
We adopt a two-phase training strategy consisting of supervised fine-tuning (SFT) followed by reinforcement fine-tuning (RFT) to stabilize event reasoning and optimize trading decisions. All experiments are conducted on $8\times$ NVIDIA A100 GPUs (40GB).
$\bullet$ Supervised fine-tuning. We first apply SFT with LoRA adaptation to initialize a reasoning-aware policy that predicts event semantics and market impact in a stable manner. The detailed hyperparameter configuration for SFT is summarized in Table 5.
$\bullet$ Reinforcement fine-tuning. Building on the initialized SFT policy, we further perform reinforcement fine-tuning using Group Relative Policy Optimization (GRPO) together with the proposed hierarchical-gated reward model, to directly optimize decision-oriented trading objectives. The detailed hyperparameter settings for GRPO are reported in Table 6.
Table 7. Summary of event-centric dataset.
Field Description
News Raw financial news article
Event time ($t_{0}$) Publication timestamp of the news
StockInfo ($i$) Associated traded asset and its basic profile
Event type ($e$) Sematic category of event
Direction ($d$) Market reaction direction: positive / negative / neutral
Strength ($s$) Trading strength indicator: strong / weak
CAR ($c$) Cumulative abnormal return
A.1.2. Dataset
We construct a large-scale Chinese equity dataset covering 5,282 A-share stocks, integrating structured market data with multi-source textual information. Formally, the dataset consists of a collection of event instances, each corresponding to a financial news event associated with a specific traded asset and timestamp. each event instance is represented as a structured record: $\mathcal{D}_{\text{event}}=\{(\text{News},t_{0},\text{StockInfo},e,d,s,c)\}$, following the format defined in Table 7. Here, News denotes the original news text, $t_{0}$ the event timestamp, StockInfo the associated traded asset, $e$ the fine-grained event type, $d$ the directional label, $s$ the strength indicator, and $c$ the event-driven cumulative abnormal return.
Daily price and volume data are obtained from Tushare, while textual signals are collected from the Datayes platform, complemented by firm-level profile information from Wind. To support stable event statistics and avoid information leakage, the dataset is chronologically partitioned into a historical window for statistical estimation, followed by non-overlapping training, validation, and test splits. Detailed data ranges and split configurations are summarized in Table 8.
To characterize temporal dependencies around each event, we adopt a standard event‑study timeline consisting of three individual intervals: an estimation window $\mathcal{T}{\text{est}}=(T_{0},T_{1}]$, an event window $\mathcal{T}{\text{evt}}=(T_{1},T_{2}]$, and a post‑event window $\mathcal{T}{\text{post}}=(T_{2},T_{3}]$, as illustrated in Figure 5. The estimation window precedes the event timestamp $t_{0}$ and provides a stable basis for estimating event‑free normal returns, ensuring that model calibration remains unaffected by upcoming market reactions. A short lag between $\mathcal{T}{\text{est}}$ and $\mathcal{T}{\text{evt}}$ is introduced to mitigate possible information leakage from early disclosures. The event window $\mathcal{T}{\text{evt}}$ captures immediate abnormal price responses, while the post‑event window $\mathcal{T}{\text{post}}$ extending to $T_{3}$ reflects subsequent return drift and adjustment dynamics. This study focuses on market reactions within $\mathcal{T}{\text{evt}}$ and does not model long‑horizon price dynamics within $\mathcal{T}{\text{post}}$, where returns generally revert toward their normal state.
Beyond scale, the dataset is distinguished by its rich semantic structure and high-quality financial annotations. Each event is categorized into a diverse set of event types, including personal behavior, equity change, asset change, dividend, risk warning, financing, financial status, violation, industry, and rating adjustment. This taxonomy is designed to cover a broad spectrum of firm-specific actions and external market signals, capturing heterogeneous information sources that are known to drive stock price movements. Moreover, each event is associated with sentiment labels and paired with its cumulative abnormal return, enabling direct supervision at both the semantic and economic levels. This dataset is intended to support future academic research on financial event understanding, market impact analysis, and event-driven modeling.
Table 8. Dataset Statistics
Country Stock Split Date Range Price Text / Company Info
CN 5282 Historical Stat 2023/01/01 – 2023/10/24 Tushare Datayes / Wind
Train 2023/10/25 – 2024/08/27
Validation 2024/08/28 – 2024/11/11
Test 2024/11/12 – 2025/01/25
Figure 5. Temporal structure of the event study framework.
A.1.3. Metrics
To evaluate the performance of Janus-Q, we adopt a set of complementary metrics that jointly assess event-level prediction accuracy, decision correctness, and trading performance.
Mean Absolute Error (MAE):
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}\left|\hat{c}_{i}-c_{i}\right|$
Root Mean Square Error (RMSE):
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\hat{c}_{i}-c_{i}\right)^{2}}$
where $\hat{c}_{i}$ is the predicted CAR, and $c_{i}$ is the true CAR. MAE measures the average deviation between the predicted and true values of abnormal returns, capturing the overall accuracy of the predictions. RMSE assesses the sensitivity of the model to large errors, penalizing large deviations more heavily than MAE.
Direction Accuracy (DA):
$\text{DA}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\text{sign}(\hat{d}_{i})=\text{sign}(d_{i})\right)$
where $\mathbb{I}$ is the indicator function, $\hat{d}_{i}$ is the predicted direction, and $d_{i}$ is the true direction. DA evaluates whether the predicted direction of the market movement aligns with the actual result achieved, providing information on the accuracy of the model’s direction prediction.
Event Type Accuracy (ETA):
$\text{ETA}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\hat{e}_{i}=e_{i}\right)$
where $\hat{e}_{i}$ is the predicted event type, and $e_{i}$ is the true event type. ETA measures the correctness of the event semantic interpretation, ensuring that the model accurately identifies the event type. Correctly distinguishing between event types enhances downstream trading performance by enabling portfolio allocations tailored to different events.
To evaluate practical trading effectiveness, we report the Sharpe Ratio (SR):
$\text{SR}=\frac{\mathbb{E}[r]}{\sigma_{r}}$
where $\mathbb{E}[r]$ is the expected return, and $\sigma_{r}$ is the standard deviation of returns. SR measures the risk-adjusted return, providing an indication of the strategy’s profitability relative to its volatility.
Maximum Drawdown (MDD):
$\text{MDD}=\max_{t}\left(\frac{\text{peak}(t)-\text{trough}(t)}{\text{peak}(t)}\right)$
where $\text{peak}(t)$ is the highest portfolio value up to time $t$, and $\text{trough}(t)$ is the lowest portfolio value after time $t$. MDD quantifies the worst peak-to-trough loss, assessing the strategy’s resilience during market downturns.
Together, these metrics provide a balanced evaluation of predictive reliability, decision consistency, and real-world trading performance.
A.2. Event empirical study
A.2.1. Historical CAR Statistics
Figure 6 illustrates the distribution of post-event cumulative abnormal returns across different event categories. We observe substantial heterogeneity in both dispersion and tail behavior among event types. Risk-related events such as Risk Warning and Violation exhibit wider distributions and heavier tails, indicating higher uncertainty and asymmetric market reactions, as well as a greater prevalence of extreme outcomes that may offer opportunities for outsized abnormal returns. In contrast, routine corporate events including Dividend and Industry announcements show more concentrated distributions around zero, suggesting relatively stable and predictable impacts. These differences highlight that event categories are associated with distinct risk return profiles, motivating event-aware modeling rather than uniform treatment of news signals.
Figure 6. Distribution of post-event cumulative abnormal returns across event categories.
A.2.2. Historical Magnitude
Figure 7 reports the mean absolute post-event CAR for each event type, reflecting the average strength of market reactions regardless of direction. Events related to regulatory or risk disclosures, such as Risk Warning and Violation, exhibit the largest magnitudes, with mean absolute CAR exceeding 0.05 and 0.03, respectively, indicating pronounced and economically meaningful market responses. In contrast, softer informational events, including Personal Behavior and Rating Adjustment, show substantially smaller impacts, with average magnitudes below 0.02. This ranking confirms that not all news carries equal economic significance and suggests that event magnitude provides a natural prior for allocating trading attention and position size, motivating magnitude-aware weighting in event-driven trading strategies. To preserve temporal relevance, such magnitude-based weights are periodically re-estimated over rolling windows, allowing the strategy to adapt to evolving market conditions and shifting event dynamics.
Figure 7. Mean absolute post-event cumulative abnormal returns by event category.
A.2.3. Impact of Event Type Weighting on Trading Performance
Figure 8. Net asset value (NAV) comparison between the Equal-Weighted and Type-Weighted Janus-Q strategies.
Figure 8 compares the net asset value (NAV) trajectories of the Equal-Weighted and Type-Weighted Janus-Q strategies. While both strategies benefit from event-driven signals, the Type-Weighted variant achieves higher cumulative returns and exhibits greater stability during periods of heightened market volatility. Notably, between early and mid-December 2024, the Equal-Weighted strategy experiences pronounced drawdowns and temporarily enters negative territory, reflecting overexposure to low-impact or noisy events. In contrast, the Type-Weighted strategy remains comparatively stable throughout the same interval, as capital allocation is concentrated on historically high-impact events. The advantage becomes most apparent around high-impact event clusters, where weighting events by their historical CAR magnitudes enables more efficient capital allocation. These results demonstrate that incorporating event-level statistical priors effectively transforms heterogeneous market reactions into tangible trading gains.
A.3. Risk Model Settings
To isolate event-driven abnormal returns from systematic risk exposures, we adopt a standard multi-factor risk model consistent with the CNE5 framework widely used in the Chinese equity market. [^4] https://www.msci.com/documents/10199/2935796a-0a80-4050-934a-12966d1e2518 CNE5 is a Barra-style equity risk model that decomposes stock returns into market, industry, and style-driven components, enabling robust neutralization of common risk premia.
In our implementation, the factor exposure vector $\mathbf{x}_{i,t}$ includes both industry factors and a set of non-financial style factors, covering: Size, Liquidity, Volatility, Momentum, and Reversal. These factors capture systematic return patterns unrelated to firm-specific events, such as liquidity shocks, short-term reversals, or broad market sentiment. We fully reproduce the factor construction and cross-sectional regression procedure following the CNE5 specification. By removing the estimated factor-driven component from market-adjusted returns, the resulting abnormal returns more accurately reflect idiosyncratic, event-induced price movements rather than generic style or industry effects.
A.4. Supplementary Experiments
A.4.1. Sensitive Analysis of Holding Period
(a) Total Return
(b) Sharpe Ratio
Figure 9. Sensitivity analysis of backtesting performance with respect to the holding period.
Based on the experimental results in Table 2, we select the three best-performing models for further evaluation. Figure 9 examines the sensitivity of performance to the holding period. As the holding horizon increases, most baseline models exhibit pronounced performance degradation. For instance, QwQ-32B achieves a positive total return of 0.0567 and a Sharpe Ratio of 1.7642 at a one-day holding period but turns negative by day three, with its Sharpe Ratio falling below 1.5-1.5 at longer horizons. A similar though milder decline is observed for DeepSeek-v3.1-nex-n1, whose total return decreases from 0.0138 at one day to 0.1169-0.1169 at ten days. This deterioration arises from the accumulation of overlapping event-driven positions, which increases exposure to unrelated market fluctuations as the holding horizon extends.
In contrast, Janus-Q demonstrates markedly greater robustness across holding periods. It reaches peak performance at short horizons, with a total return of 0.122 and a Sharpe Ratio of 1.8074 at one day, while maintaining positive returns up to nine days. Even at a ten-day horizon, its performance declines smoothly rather than collapsing, retaining a positive Sharpe Ratio of 0.398. This stability suggests that Janus-Q implicitly models the temporal decay of event influence and avoids excessive position persistence.
The CSI 300 benchmark remains unchanged across horizons, as it follows a static buy-and-hold strategy that is insensitive to event timing. Overall, these findings indicate that in unconstrained settings, event-driven strategies benefit from shorter holding periods that align with the transient nature of news impacts. Prolonged horizons lead to overlapping exposures and amplified drawdowns, while appropriate position control and horizon-aware exposure limits, examined later in our experiments, can help mitigate such degradation.
(a) Holding Period = 5
(b) Holding Period = 10
Figure 10. Sensitivity analysis of backtesting performance with respect to the maximum position ratio.
A.4.2. Sensitive Analysis of Maximum Position Ratio
Building on the previous analysis of holding-period sensitivity, Figure 10 presents a complementary experiment examining how position control affects performance under different trading frequencies. Specifically, it analyzes backtesting performance with respect to the maximum position ratio, using 5-day and 10-day holding periods as representative examples.
The maximum position ratio constrains the portfolio’s total notional exposure relative to its net asset value (NAV). A ratio of $k\times$ limits the combined exposure of all open positions to $k$ times the current NAV, thereby regulating both leverage and the extent to which multiple signals can be executed concurrently. The $\infty$ setting corresponds to the default configuration in the main experiments, where no explicit upper bound on the position ratio is applied.
As shown in Figure 10, excessively restrictive position limits (e.g., $1\times$) tend to impair performance, as the strategy becomes capital-constrained and must strictly adhere to the fixed holding period, missing profitable intermediate signals. Increasing the limit to moderate levels (e.g., $2\times$ or $3\times$) generally enhances performance by enabling partial overlap among positions and fuller utilization of concurrent signals. However, relaxing the constraint too much can again degrade performance, as excessive exposure leads to the accumulation of stale positions whose diminishing returns offset gains from newly arriving events.
Overall, these results reveal a trade-off between capital efficiency and signal freshness, indicating that moderate position limits provide the most balanced performance. Moreover, appropriate position control allows event-driven strategies to adapt to longer holding horizons and different trading frequencies, effectively stabilizing returns while maintaining responsiveness to new information.
Algorithm 1 Hierarchical-Gated Reward Modeling (HGRM)
1:Model response $\mathcal{R}$, ground truth $(c,e)$, threshold $\tau$, transaction cost $\kappa$, clip bound $\rho$, discount $\alpha$
2:Reward $R$
3:Parse prediction $\{\hat{c},\hat{d},\hat{s},\hat{e}\}$ from $\mathcal{R}$
4:if $\hat{d}$ is missing then
[…]
6:end if
7:if $\hat{s}$ is missing then
8:  $\hat{s}\leftarrow\texttt{strong}$ if $|\hat{c}|>\tau$ else weak
9:end if
10:Derive $d\leftarrow\mathrm{sign}(c)$, $s\leftarrow\texttt{strong}$ if $|c|>\tau$ else weak
11:// Hard gate: direction correctness
12:Compute direction score $s_{\mathrm{dir}}(\hat{d},d)$
13:Set hard gate $g_{\mathrm{dir}}\leftarrow 1(s_{\mathrm{dir}}\geq 0)$
14:// Soft modulation: event-type consistency
15:Compute event score $s_{\mathrm{evt}}(\hat{e},e)$
16:Set discount factor $m_{\mathrm{evt}}\leftarrow 1$ if $\hat{e}=e$ else $\alpha$
17:// Cost-aware trading payoff
18:Compute $\mathrm{pnl}\leftarrow\mathrm{PnL}(\hat{d},c,\kappa)$
19:if $g_{\mathrm{dir}}=0$ then
20:  $r_{\mathrm{pnl}}\leftarrow 0$ $\triangleright$ no trade executed
21:else
22:  $r_{\mathrm{pnl}}\leftarrow\mathrm{clip}(m_{\mathrm{evt}}\cdot\mathrm{Pnl},-\rho,\rho)$
23:end if
24:// Magnitude shaping and process reward
25:if $g_{\mathrm{dir}}=1$ then
26:  $r_{\mathrm{mag}}\leftarrow\exp(-|\hat{c}-c|/\sigma)$
27:  Compute process reward $r_{\mathrm{proc}}$
28:else
29:  $r_{\mathrm{mag}}\leftarrow 0$; $r_{\mathrm{proc}}\leftarrow 0$
30:end if
31:// Final hierarchical reward
[…]
33:return $R$
A.5. Prompts in Janus-Q
A.5.1. Prompt for Reasoning Chain of Thought
Figure 11 illustrates the prompt used to elicit structured reasoning chains for post-hoc analysis of event-driven market reactions. The prompt guides the model to explain event categorization and observed cumulative abnormal returns by explicitly considering event characteristics, market expectations, and historical context. By enforcing a fixed multi-part structure, this prompt facilitates interpretable reasoning and supports qualitative inspection of the model’s event-to-market understanding.
A.5.2. Prompt for Training Template
Figure 12 presents the training prompt template used during supervised fine-tuning. The template requires the model to jointly predict event type, impact direction, trading intensity, and expected abnormal return, while producing a concise rationale. Historical CAR statistics are provided as contextual references rather than hard constraints, encouraging the model to balance empirical priors with event-specific reasoning. This design aligns model outputs with downstream decision objectives in event-driven trading.
Figure 11. Prompt for reasoning chain of thought.
Figure 12. Prompt for training template.